OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (aonz):

what do you think it is?

OpenStudy (anonymous):

tbh i used to understand this but i kinda forgot now :( i know it has something to do with -y and -x but im not sure

OpenStudy (aonz):

so the radius is 8^2 right?

OpenStudy (anonymous):

yes

OpenStudy (aonz):

so what do you think it is now?

OpenStudy (anonymous):

is there any formula that you could give me? im honestly really lost :c sorry

OpenStudy (aonz):

(x -h)^2 + (y - k)^2 =r^2

OpenStudy (anonymous):

ok so (2-h)^2+(1=k)^2=8^2 ??

OpenStudy (anonymous):

(1+k)***

OpenStudy (aonz):

oh umm h is 2 and k is -1

OpenStudy (anonymous):

ohhh right! (x-2)^2+(y-1)^2=8^2 ohhh ok xD thank you

OpenStudy (aonz):

its y+1

OpenStudy (anonymous):

oh right! cuz negative turns to positive >.< sorry i always forget that! thank you though!!

OpenStudy (aonz):

No problem :)